Find an orthonormal basis of R^3 of which (1√10,0,-3/√10) is one element

1
Expert's answer
2021-12-05T18:29:29-0500

The length of the vector "\vec v_1" is "1" as


"||\vec v_1 ||=\sqrt{(1/\sqrt{10})^2+(0)^2+(-3/\sqrt{10})^2}=1"

We want to find two vectors "\vec v_2,\vec v_3" such that "\{\vec v_1,\vec v_2,\vec v_3\}"is an orthonormal basis for "\R^3."

Let "\vec u=\langle x, y, z \rangle" be a vector that is perpendicular to "\vec v_1."

Then we have "\vec u\cdot \vec v_1=0," and hence we have the relation


"(1/\sqrt{10})(x)+(0)(y)+(-3/\sqrt{10})(z)=0"

"x-3z=0"

For example, the vector "\vec u=\langle 3, 1, 1 \rangle" satisfies the relation, anf hence "\vec u\cdot \vec v_1=0."

Let us define the third vector "\vec w" to be the cross product of "\vec u" and "\vec v_1"


"\vec w=\vec u\times\vec v_1=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 3 & 1 & 1 \\
 1/\sqrt{10} & 0 & -3/\sqrt{10}
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & 1 \\
 0 & -3/\sqrt{10}
\end{vmatrix}-\vec j\begin{vmatrix}
 3 & 1 \\
 1/\sqrt{10} & -3/\sqrt{10}
\end{vmatrix}"

"+\vec k\begin{vmatrix}
 3 & 1 \\
 1/\sqrt{10} & 0
\end{vmatrix}"


"=(-3/\sqrt{10})\vec i+(10/\sqrt{10})\vec j+(-1/\sqrt{10})\vec k"

By the property of the cross product, the vector "\vec w" is perpendicular to both "\vec u, \vec v_1."

"||\vec u||=\sqrt{(3)^2+(1)^2+(1)^2}=\sqrt{11}"

"||\vec w||=\sqrt{(-3/\sqrt{10})^2+(10/\sqrt{10})^2+(-1/\sqrt{10})^2}"

"=\sqrt{11}"

Therefore, the set


"\{\vec v_1,\vec v_2,\vec v_3\}"


"=\bigg\{\begin{bmatrix}
 1/\sqrt{10} \\
 0 \\
-3/\sqrt{10}
\end{bmatrix},\begin{bmatrix}
 3/\sqrt{11} \\
 1/\sqrt{11}\\
1/\sqrt{11}
\end{bmatrix},\begin{bmatrix}
 -3/\sqrt{110} \\
 10/\sqrt{110}\\
-1/\sqrt{110}
\end{bmatrix}\bigg\}"

is an orthonormal basis for "\R^3."